if x1 and x2 complementary angles and if the measure if x1 is 18 degrees more than the measure of x2, determine the measures of x1 and x2
Answer by stanbon(75887)

if x1 and x2 complementary angles and if the measure if x1 is 18 degrees more than the measure of x2, determine the measures of x1 and x2
-------------
Equations:
x1 + x2 = 90
x1 = x2 + 18
------------------
Substitute for x1 and solve for x2:
x2+18 + x2 = 90
2*x2 = 72
x2 = 36 degrees
x1 = 36+18 = 54 degrees
============================
Cheers,
Stan H.